40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1187     Entry time: Mon Dec 8 11:54:27 2008
Author: Yoichi 
Type: Update 
Category: General 
Subject: IFO mirrors aligned 
This morning, I re-aligned the IFO mirrors to see if they were badly moved by the earthquake.
The both arms locked just by the restoring scripts, but the transmission was about 0.7. So I aligned them
with the dithering scripts.
To lock the PRMI, I had to manually tweak the PRM alignment. After running the dithering script, the SPOB
went up to 1200.
I also had to tweak the SRM to get the DRMI locked. After the dithering script, the SPOB was 4200 and REFL166Q
was 3000.
ELOG V3.1.3-